A rectangular park is 30 miles long and 10 miles wide. How long is the pedestrian route that runs diagonally across the park?
Answer by mananth(16949)

A rectangular park is 30 miles long and 10 miles wide. How long is the pedestrian route that runs diagonally across the park?
Use pythagoras theorem to find the diagonal
30^2+10^2=D^2
900+100=D^2
1000=D^2
D= 31.6 miles the diagonal distance